OpenStudy (anonymous):

How would I use Ohm's Law to calculate the current (in amps) traveling through a 100 ohm resistors in series with a 3 V battery? I would like to compare this current to the current traveling through a 1 kOhm resistor as well. My next question is, once I know the value for current above, how much would I have to increase the voltage using the 1 kOhm resistor to create a current equal to the value for current above? Finally, if I had a lamp that used 3W of power when connected to a 6 V battery, how would I calculate the internal resistance of the lamp and current flowing through filament?

OpenStudy (anonymous):

From ohm's law, V=IR. Therefore, 3=100*I, I=0.03 amps. In the second question, V=1000*0.03, V=30V. In the third part, power, P=V*I. therefore, 3=6*I. find out I, use ohms law again and find it!